A hotel has 200 rooms. Those with kitchen facilities rent for $100 per night and those without kitchen facilities rent for $80 per night. On a night when the hotel was completely occupied, revenues were $17,000. How many of each type of room does the hotel have?
Answer by mananth(16946)

rooms with kitchen x
rooms without kitchen ------------ y
x+y=200 .............1
100x+80 y=17000 .............2
Eliminate y
multiply (1)by 80
Multiply (2) by -1
80x+5y=16000
-100x+2 y=-17000
Add the two equations
-20x=-1000
/-20
x=50 rooms with kitchen

plug value of x in (1)

x+y=200
50 + y=200
y=150 rooms without kitchen